# Leveraging Symmetry Merging in Pauli Propagation

### v1

@@ I Introduction
Understanding the dynamics of quantum operators lies at the heart of quantum many-body physics and quantum information theory. Among many practical applications, operator dynamics provide physical insights into quantum information scrambling, many-body chaos, and dynamical correlations in interacting systems for quantum metrology and sensing [plodzien2025lecture, sieberer2019digital, nandy2025quantum, montenegro2025quantum].
There now exists a broad zoo of classical simulation methods for approximating quantum dynamics. Matrix product states (MPS) have been remarkably successful in one dimension [Paeckel_2019, Haegeman_2011, Vidal_2004, Verstraete_2004, White_2004], and a variety of higher-dimensional generalizations are advancing rapidly [mandra2025heuristic, Alkabetz_2021, Tindall_2023, tindall2025dynamics, rudolph2025simulating, gray2025tensor]. Additional theoretical frameworks include neural quantum states [Carleo_2017, carleo2024simulatingadiabaticquantumcomputation], mean-field approaches to operator hydrodynamics [khemaniOperatorSpreadingEmergence2018, rakovszkyDiffusiveHydrodynamicsOutTimeOrdered2018], and specialized semiclassical techniques such as the discrete truncated Wigner approximation [wurtz2018cluster]. Yet no single “winner-takes-all” method exists that reliably handles all problems.
@@
In many scenarios, it is effective to study the evolution of observables in the Heisenberg picture rather than tracking the full state evolution. This idea has been explored using MPS to simulate operators dynamics in one dimension via dissipation-assisted operator evolution (DAOE) [rakovszkyDissipationassistedOperatorEvolution2022]. Another approach to approximate operator dynamics is through Pauli propagation (PP), which tracks the coefficients of Pauli operators under successive circuit layers [rallSimulationQubitQuantum2019, aharonov2022polynomial, beguvsic2023simulating, fontana2023classical, shao2023simulating, rudolph2023classical, schuster2024polynomial, angrisani2024classically, gonzalez2024pauli, lerch2024efficient, cirstoiu2024fourier, angrisani2025simulating, fuller2025improved, rudolph2025simulating, angrisani2025simulating, rakovszkyDissipationassistedOperatorEvolution2022]. Pauli propagation encodes Pauli strings as bitstrings and provides a fast bitwise processing of evolution for dynamical quantities involving two procedures: 1) truncating irrelevant Pauli operators and 2) merging Paulis that are identical [rudolph2025pauli]. However, as the system evolves, the number of nonzero Pauli coefficients typically grows exponentially, leading to severe memory and runtime bottlenecks that limit standard Pauli propagation [loizeauQuantumManybodySimulations2024, dowling2025bridging, dowling2025magic].
@@
Our symmetry-merging Pauli propagation (symmetry PP) algorithm generalizes naturally across any discrete symmetry groups and can be combined with noise models or other truncation schemes, as illustrated in our numerical examples. The resulting method achieves both improved stability and reduced space complexity, while memory is a key limiting factor of the standard Pauli propagation (standard PP).
@@
Consider a unitary $U$ composed of $L$ layers $U=\prod_{l=1}^{L}U_{l}$. For instance, the $l$-th unitary layer $U_{l}$ can be a trotterized Hamiltonian layer $U_{l}=\prod_{j}e^{-ih^{(1)}_{j}t}e^{-ih^{(2)}_{j}t}$ for the translational invariant Hamiltonian $H=\sum_{j}h^{(1)}_{j}+h^{(2)}_{j}$ or only a sublayer satisfying the symmetry $U_{l}=\prod_{j}e^{-ih^{(1)}_{j}t}$. Let $G$ be a finite symmetry group and $A_{g}$ be the representation of $g\in G$. Each layer is invariant under $G$ defined by the group action $g\cdot(\,)$:
@@
Under the evolution of a symmetric unitary, this redundancy allows observables to be computed from a reduced set of representative $n$-qubit Pauli strings, denoted as $\mathcal{R}^{G}_{n}$. While we focus on unitary dynamics, our framework generalizes for open quantum systems. We illustrate this flexibility in our numerical example by adding noise layers modeling a local Lindbladian dynamics [lidar2019lecture, breuer2002theory].
@@ II Symmetry-merging Pauli propagation
Given an initial state $\rho$ and the evolution unitary $U=\prod_{l=1}^{L}U_{l}$ with discrete symmetry $G$, our goal is to compute the expectation value of the evolved target observable $\mathrm{Tr}[\rho\,U^{\dagger}OU]$. At a high level, the standard PP algorithm evolves a target observable $O$ using the Heisenberg equation of motion. The evolved operator expectation value can be expanded in the Pauli basis leading to a Pauli path integral,
@@ IV Numerical examples
where $\langle ij\rangle$ enumerates the nearest-neighbor sites and $h_{x}(h_{z})$ is the transverse (longitudinal) field. It is known that in the absence of the longitudinal field, the Hamiltonian is integrable using Jordan Wigner transformations [lieb1961two] such that only a subspace is explored, which is spanned by a polynomial in $n$ number of Pauli operators. Away from the integrable limit ($h_{z}=0$), this model has been widely used to study thermalization and ergodic dynamics [hertz2018quantum, roberts2015localized]. In our numerical results, we consider the field strengths $g_{x}=1.4,g_{z}=0.9045$ studied in prior works [rakovszkyDissipationassistedOperatorEvolution2022]. Under periodic boundary conditions, the system has the translation symmetry $\mathbb{Z}_{n}$.
@@
Example: all-to-all XXZ Heisenberg model. To demonstrate how symmetry PP allows us to simulate physically relevant dynamics, we investigate the dynamics of the all-to-all interacting Heisenberg model with long-range power law interactions decaying with a distance defined on a $2$-dimensional lattice over an exponent $\alpha$. This model is a rich playground, as the ground state phase diagrams exhibit a complex structure in entanglement and critical phenomena [frerot2017entanglement]. Moreover, it is also relevant for quantum simulations and metrology in atom-based platforms such as the Rydberg atom arrays, dipolar molecules and trapped ion arrays [perlin2020spin, bilitewski2021dynamical, muleady2023validating]. The power-law XXZ Hamiltonian is given by
@@
The unitary is the trotterized Hamiltonian evolution $U=\prod_{l=1}^{L}\prod_{i<j}e^{-ih_{ij}\delta}$ where the time step is $\delta=t/L$ and $h_{ij}=-J_{\perp}\left(X_{i}X_{j}+Y_{i}Y_{j}+(\Delta+1)Z_{i}Z_{j}\right)/|\mathbf{r}_{i}-\mathbf{r}_{j}|^{\alpha}$ is the local Hamiltonian operator. It is possible to merge after only evolving with sublayers to further improve the efficiency. Because pairs of Pauli operators commute with each other e.g. $\left[X_{i}X_{j},\,Y_{i}Y_{j}\right]=0$, we can further decompose the rotation into Pauli rotations as $e^{i\left(\theta_{x}X_{i}X_{j}+\theta_{y}Y_{i}Y_{j}+\theta_{z}Z_{i}Z_{j}\right)}=e^{i\theta_{x}X_{i}X_{j}}e^{i\theta_{y}Y_{i}Y_{j}}e^{i\theta_{z}Z_{i}Z_{j}}$. Alternative strategies include higher order trotterization with multi product formula [childs2021theory], finding a minimum set of grouped operators [berry2007efficient] and randomized trotterization [campbell2019random, childs2019faster]. In the present example, we have numerically observed that the trotterization order affects the relevant observable expectation values minimally.
We will study the total spin operator [muleady2023validating]
@@
where $\sigma_{\mu,i}$ are spin-half Pauli matrices (e.g. $\sigma_{x,i}=\frac{1}{2}X_{i}$). The measurement operator and unitary evolution are translationally invariant, therefore we apply our merging algorithm with translation symmetries along both $x$ and $y$ directions independently. While there exists other specialized methods such as discrete truncated Wigner approximation [wurtz2018cluster], this dynamics is highly non-trivial for MPS-based simulations [muleady2023validating] owing to the all-to-all connected circuit topology. In our naive MPO simulations, we observe difficulty in convergence for simulation time $t\bar{J}\lvert\Delta\rvert>0.02$ using a bond dimension up to $400$.
In Fig. 3, we demonstrate that symmetry PP allows us to simulate the dynamics more accurately than standard PP. To simulate effects of presence of noise, our numerical simulations include a layer of depolarizing noise with strength $\gamma$ after each trotterization step. The noise layer can be viewed as a time discretized model for a local Lindbladian [breuer2002theory, lidar2019lecture]. In addition, we adopt a coefficient truncation strategy such that only terms with coefficients greater than a truncation threshold $\epsilon$ are kept [rudolph2025pauli]. For instance, a truncated operator is given by $O^{\text{trunc}}=\sum_{\alpha\,|\,c_{\alpha}\leq\epsilon}c_{\alpha}s_{\alpha}$ for some small $\epsilon$. We study the effect of different truncation thresholds on simulation accuracy and find that symmetry PP achieves faster convergence even for a larger truncation threshold $\epsilon=10^{-3}$ (circles) compared to a truncation threshold at $\epsilon=5\times 10^{-5}$ (solid line). In comparison, standard PP diverges at earlier times (dashed vs. solid lines). Moreover, symmetry PP with a higher truncation threshold $\epsilon=0.05$ agrees with standard PP at a much lower threshold ($\epsilon=0.001$). This can be understood by realizing that symmetry merging tends to increase the magnitudes of the coefficients, in turn increasing the Pauli strings that are kept per threshold. While it is inevitably exponentially costly to simulate long-time dynamics, given a fixed budget of memory and computation time, symmetry PP converges to longer time and leads to more accurate results hence reducing the computational resources for a target accuracy.
@@ V Discussion
We have established a group-theoretic symmetry-merging framework for Pauli propagation, called symmetry PP. By propagating only a set of Pauli orbit representatives, we reduce the memory cost by a factor related to the size of the group. Such a memory saving is highly impactful for future improvements using graphics processing unit (GPU) because the acceleration hardwares typically have more constraints in memory. Most recently, related works have also taken symmetry into consideration via merging in PP [loizeauQuantumManybodySimulations2024, haghshenas2025digital]. Theoretically, we provide rigorous guarantees for the correctness and computational complexity. Then we demonstrate the effectiveness of symmetry PP for an ergodic dynamics of $1d$ tilted-field Ising model and an all-to-all Heisenberg dynamics. We observe an improved stability in symmetry PP simulations, despite subtleties in the memory saving due to an unavoidable exponentially large space of Pauli strings. It would be interesting to analyze the improvement by considering magic monotones in resource theory [dowling2025magic, turkeshi2025magic].
For future work, it would be interesting to explore symmetric merging strategies in other propagation based methods such as Majorana propagation for fermionic systems [miller2025simulation, alam2025fermionic, alam2025programmable, danna2025majorana] or displacement propagation for bosonic system [upreti2025quantum]. This includes any future merging strategies for other discrete symmetries. More ambitiously, developing propagation subroutines for continuous symmetries will be highly relevant and may yield substantial reductions in computational cost. An abstract analysis of this direction has already been initiated using representation theory in Ref. [cirstoiu2024fourier]. Such progress would open new avenues for investigating the role of symmetries in physically relevant quantum Floquet dynamics, such as those previously explored in random circuits with $U(1)$ symmetry [khemaniOperatorSpreadingEmergence2018, rakovszkyDiffusiveHydrodynamicsOutTimeOrdered2018].
Note: While preparing this manuscript we used similar methods with our collaborators in Ref. [haghshenas2025digital] and a related approach was used in the context of computing Lanczos coefficients and the Runge-Kutta method in Ref. [loizeauQuantumManybodySimulations2024]. Additionally, symmetry in Krylov methods has recently been considered in Ref. [loizeau2025opening].
@@
Acknowledgments. The authors acknowledge Nathan Leitao for insightful discussions. YT acknowledges support from NCCR spin, a National Centre of Competence in Research, funded by the Swiss National Science Foundation (grant number 565785). SYC was supported by Laboratory Directed Research and Development (LDRD) program of Los Alamos National Laboratory (LANL) under project number 20260043DR and by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research under Contract No. DE-AC05-00OR22725 through the Accelerated Research in Quantum Computing Program MACH-Q project. MSR acknowledges funding from the 2024 Google PhD Fellowship and the Swiss National Science Foundation [grant number 200021-219329]. ZH acknowledges support from the Sandoz Family Foundation-Monique de Meuron program for Academic Promotion.
[…]
oneΔ
@@ A Pauli Path Methods
In this appendix, we review the framework of Pauli path based simulations, adopting the conventions of Pauli propagation [rudolph2025pauli]. Let us consider an approximately maximally mixed state, commonly studied in operator spreading,
@@
Here, the operators live in the $n$-qubit Hilbert space $\mathcal{H}=(\mathbb{C}^{2})^{\otimes n}$, with an inner product $\langle A,B\rangle=\mathrm{Tr}[A^{\dagger}B]$. The operators can be represented by matrices $M_{2^{n}}(\mathbb{C})$. To analyze quantum dynamics, it is useful to vectorize these operators in an enlarged Hilbert space, called the Liouville space $\mathcal{L}=\mathcal{H}\otimes\mathcal{H}^{*}$ [gyamfi2020fundamentals]. In the vectorization formalism, an operator $A$ on $\mathcal{H}$ can be uniquely mapped to a vector $\mbox{$|A\rangle\rangle$}\in\mathcal{L}$ as the following:

### v2

@@ I Introduction
Understanding the dynamics of quantum operators lies at the heart of quantum many-body physics and quantum information theory. Among many practical applications, operator dynamics provide physical insights into quantum information scrambling, many-body chaos, and dynamical correlations in interacting systems for quantum metrology and sensing [45, 58, 41, 39].
There now exists a broad zoo of classical simulation methods for approximating quantum dynamics. Matrix product states (MPS) have been remarkably successful in one dimension [42, 28, 64, 63, 65], and a variety of higher-dimensional generalizations are advancing rapidly [37, 4, 59, 60, 54, 26]. Additional theoretical frameworks include neural quantum states [14, 13], mean-field approaches to operator hydrodynamics [31, 47], and specialized semiclassical techniques such as the discrete truncated Wigner approximation [66]. Yet no single “winner-takes-all” method exists that reliably handles all problems.
@@
In many scenarios, it is effective to study the evolution of observables in the Heisenberg picture rather than tracking the full state evolution. This idea has been explored using MPS to simulate operators dynamics in one dimension via dissipation-assisted operator evolution (DAOE) [48]. Another approach to approximate operator dynamics is through Pauli propagation (PP), which tracks the coefficients of Pauli operators under successive circuit layers [49, 1, 8, 22, 57, 52, 56, 6, 25, 32, 17, 5, 24, 54, 5, 48]. Pauli propagation encodes Pauli strings as bitstrings and provides a fast bitwise processing of evolution for dynamical quantities involving two procedures: 1) truncating irrelevant Pauli operators and 2) merging Paulis that are identical [53]. However, as the system evolves, the number of nonzero Pauli coefficients typically grows exponentially, leading to severe memory and runtime bottlenecks that limit standard Pauli propagation [36, 21, 20].
@@
Our symmetry-merging Pauli propagation (symmetry PP) algorithm generalizes naturally across any discrete symmetry groups and can be combined with noise models or other truncation schemes, as illustrated in our numerical examples. The resulting method achieves both improved stability and reduced space complexity, while memory is a key limiting factor of the standard Pauli propagation (standard PP). Various works have exploited symmetries in one way or another [36, 35, 19, 29, 43]. We present the first principled Pauli propagation algorithm with exhaustive symmetry merging throughout the simulation and with principled guarantees for correctness and performance.
@@
Consider a unitary $U$ composed of $L$ layers $U=\prod_{l=1}^{L}U_{l}$. For instance, the $l$-th unitary layer $U_{l}$ can be a trotterized Hamiltonian layer $U_{l}=\prod_{j}e^{-ih^{(1)}_{j}t}e^{-ih^{(2)}_{j}t}$ for the translational invariant Hamiltonian $H=\sum_{j}h^{(1)}_{j}+h^{(2)}_{j}$ or only a sublayer $U_{l}=\prod_{j}e^{-ih^{(1)}_{j}t}$. Let $G$ be a finite symmetry group and $A_{g}$ be the representation of $g\in G$. Each layer is invariant under $G$ defined by the group action $g\cdot(\,)$:
@@
Under the evolution of a symmetric unitary, this redundancy allows observables to be computed from a reduced set of representative $n$-qubit Pauli strings, denoted as $\mathcal{R}^{G}_{n}$. While we focus on unitary dynamics, our framework generalizes for open quantum systems. We illustrate this flexibility in our numerical example by adding noise layers modeling a local Lindbladian dynamics [33, 11].
@@ II Symmetry-merging Pauli propagation
Given an initial state $\rho$ and the unitary $U=\prod_{l=1}^{L}U_{l}$ with discrete symmetry $G$, our goal is to compute the expectation value of the evolved target observable $\mathrm{Tr}[\rho\,U^{\dagger}OU]$. At a high level, the standard PP algorithm evolves a target observable $O$ using the Heisenberg equation of motion. The evolved operator can be expanded in the Pauli basis as a Pauli path integral,
@@ IV Numerical examples
where $\langle ij\rangle$ enumerates the nearest-neighbor sites and $h_{x}(h_{z})$ is the transverse (longitudinal) field. It is known that in the absence of the longitudinal field, the Hamiltonian is integrable using Jordan Wigner transformations [34] such that only a subspace is explored, which is spanned by a polynomial in $n$ number of Pauli operators. Away from the integrable limit ($h_{z}=0$), this model has been widely used to study thermalization and ergodic dynamics [30, 50]. In our numerical results, we consider the field strengths $g_{x}=1.4,g_{z}=0.9045$ studied in prior works [48]. Under periodic boundary conditions, the system has the translation symmetry $\mathbb{Z}_{n}$.
@@
Example: all-to-all XXZ Heisenberg model. To demonstrate how symmetry PP allows us to simulate physically relevant dynamics, we investigate the dynamics of the all-to-all interacting Heisenberg model with long-range power law interactions decaying with a distance defined on a $2$-dimensional lattice over an exponent $\alpha$. This model is a rich playground, as the ground state phase diagrams exhibit a complex structure in entanglement and critical phenomena [23]. Moreover, it is also relevant for quantum simulations and metrology in atom-based platforms such as the Rydberg atom arrays, dipolar molecules and trapped ion arrays [44, 10, 40]. The power-law XXZ Hamiltonian is given by
@@
The unitary is the trotterized Hamiltonian evolution $U=\prod_{l=1}^{L}\prod_{i<j}e^{-ih_{ij}\delta}$ where the time step is $\delta=t/L$ and $h_{ij}=-J_{\perp}\left(X_{i}X_{j}+Y_{i}Y_{j}+(\Delta+1)Z_{i}Z_{j}\right)/|\mathbf{r}_{i}-\mathbf{r}_{j}|^{\alpha}$ is the local Hamiltonian operator. It is possible to merge after only evolving with sublayers to further improve the efficiency. Because pairs of Pauli operators commute with each other e.g. $\left[X_{i}X_{j},\,Y_{i}Y_{j}\right]=0$, we can further decompose the rotation into Pauli rotations as $e^{i\left(\theta_{x}X_{i}X_{j}+\theta_{y}Y_{i}Y_{j}+\theta_{z}Z_{i}Z_{j}\right)}=e^{i\theta_{x}X_{i}X_{j}}e^{i\theta_{y}Y_{i}Y_{j}}e^{i\theta_{z}Z_{i}Z_{j}}$. Alternative strategies include higher order trotterization with multi product formula [16], finding a minimum set of grouped operators [9] and randomized trotterization [12, 15]. In the present example, we have numerically observed that the trotterization order affects the relevant observable expectation values minimally.
We will study the total spin operator [40]
@@
where $\sigma_{\mu,i}$ are spin-half Pauli matrices (e.g. $\sigma_{x,i}=\frac{1}{2}X_{i}$). The measurement operator and unitary evolution are translationally invariant, therefore we apply our merging algorithm with translation symmetries along both $x$ and $y$ directions independently. While there exists other specialized methods such as discrete truncated Wigner approximation [66], this dynamics is highly non-trivial for MPS-based simulations [40] owing to the all-to-all connected circuit topology. In our naive MPO simulations, we observe difficulty in convergence for simulation time $t\bar{J}\lvert\Delta\rvert>0.02$ using a bond dimension up to $400$.
In Fig. 3, we demonstrate that symmetry PP allows us to simulate the dynamics more accurately than standard PP. To simulate effects of presence of noise, our numerical simulations include a layer of depolarizing noise with strength $\gamma$ after each trotterization step. The noise layer can be viewed as a time discretized model for a local Lindbladian [11, 33]. In addition, we adopt a coefficient truncation strategy such that only terms with coefficients greater than a truncation threshold $\epsilon$ are kept [53]. For instance, a truncated operator is given by $O^{\text{trunc}}=\sum_{\alpha\,|\,c_{\alpha}\leq\epsilon}c_{\alpha}s_{\alpha}$ for some small $\epsilon$. We study the effect of different truncation thresholds on simulation accuracy and find that symmetry PP achieves faster convergence even for a larger truncation threshold $\epsilon=10^{-3}$ (circles) compared to a truncation threshold at $\epsilon=5\times 10^{-5}$ (solid line). In comparison, standard PP diverges at earlier times (dashed vs. solid lines). Moreover, symmetry PP with a higher truncation threshold $\epsilon=0.05$ agrees with standard PP at a much lower threshold ($\epsilon=0.001$). This can be understood by realizing that symmetry merging tends to increase the magnitudes of the coefficients, in turn increasing the Pauli strings that are kept per threshold. While it is inevitably exponentially costly to simulate long-time dynamics, given a fixed budget of memory and computation time, symmetry PP converges to longer time and leads to more accurate results hence reducing the computational resources for a target accuracy.
@@ V Discussion
We have established a group-theoretic symmetry-merging framework for Pauli propagation, called symmetry PP. By propagating only a set of Pauli orbit representatives, we reduce the memory cost by a factor related to the size of the group. Such a memory saving is highly impactful for future improvements using graphics processing unit (GPU) because the acceleration hardwares typically have more constraints in memory. Most recently, related works have also taken symmetry into consideration via merging in PP [36, 29]. Theoretically, we provide rigorous guarantees for the correctness and computational complexity. Then we demonstrate the effectiveness of symmetry PP for an ergodic dynamics of $1d$ tilted-field Ising model and an all-to-all Heisenberg dynamics. We observe an improved stability in symmetry PP simulations, despite subtleties in the memory saving due to an unavoidable exponentially large space of Pauli strings. It would be interesting to analyze the improvement by considering magic monotones in resource theory [20, 61].
For future work, it would be interesting to explore symmetric merging strategies in other propagation based methods such as Majorana propagation for fermionic systems [38, 2, 3, 18] or displacement propagation for bosonic system [62]. This includes any future merging strategies for other discrete symmetries. More ambitiously, developing propagation subroutines for continuous symmetries will be highly relevant and may yield substantial reductions in computational cost. An abstract analysis of this direction has already been initiated using representation theory in Ref. [17]. Such progress would open new avenues for investigating the role of symmetries in physically relevant quantum Floquet dynamics, such as those previously explored in random circuits with $U(1)$ symmetry [31, 47].
Note: While preparing this manuscript we used similar methods with our collaborators in Ref. [29].
@@
Acknowledgments. The authors acknowledge Nathan Leitao for insightful discussions. YT acknowledges support from NCCR spin, a National Centre of Competence in Research, funded by the Swiss National Science Foundation (grant number 565785). SYC was supported by Laboratory Directed Research and Development (LDRD) program of Los Alamos National Laboratory (LANL) under project number 20260043DR and by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research under Contract No. DE-AC05-00OR22725 through the Accelerated Research in Quantum Computing Program MACH-Q project. MSR acknowledges funding from the 2024 Google PhD Fellowship and the Swiss National Science Foundation [grant number 200021-219329]. ZH acknowledges support from the Sandoz Family Foundation-Monique de Meuron program for Academic Promotion. We acknowledge support of the NCCR MARVEL, a National Centre of Competence in Research, funded by the Swiss National Science Foundation (grant number 205602).
@@ A Pauli Path Methods
In this appendix, we review the framework of Pauli path based simulations, adopting the conventions of Pauli propagation [53]. Let us consider an approximately maximally mixed state, commonly studied in operator spreading,
@@
Here, the operators live in the $n$-qubit Hilbert space $\mathcal{H}=(\mathbb{C}^{2})^{\otimes n}$, with an inner product $\langle A,B\rangle=\mathrm{Tr}[A^{\dagger}B]$. The operators can be represented by matrices $M_{2^{n}}(\mathbb{C})$. To analyze quantum dynamics, it is useful to vectorize these operators in an enlarged Hilbert space, called the Liouville space $\mathcal{L}=\mathcal{H}\otimes\mathcal{H}^{*}$ [27]. In the vectorization formalism, an operator $A$ on $\mathcal{H}$ can be uniquely mapped to a vector $\mbox{$|A\rangle\rangle$}\in\mathcal{L}$ as the following: